# Initial Guessing Bias: How Untrained Networks Favor Some Classes

### v4

@@ F.1 Non-null mean activation function
Let us consider a generic activation function $g\left({\cdot}\right)$, such that
$\left\langle{g^{(1)}_{i}}\right\rangle\equiv\left\langle{g\left({h^{(1)}_{i}}\right)}\right\rangle\stackrel{{\scriptstyle}}{{=}}\left\langle{g^{(1)}}\right\rangle\neq 0.$ (124)
where the second step follow from the fact that $\{h^{(1)}_{i}\}$ are identically distributed. Starting from this hypothesis on the activation function (Eq. (124)) we can follow exactly the same analysis presented in App. G.3.1. In particular:
@@
$p_{O^{(c)}}^{(\chi)}\left({x}\right)=p_{h^{(2)}_{c}}^{(\chi)}\left({x}\right)\,.$ (126)
For a generic deep architecture we can keep following the analysis of App. G.3.1, substituiting the right expression to $p_{g^{(2)}_{i}}^{(\chi)}\left({x}\right)$ dependent to the specific used activation function. Proceeding in this way we will find, for each layer, a symmetry breaking in the nodes distribution, similarly to what we observed in the second layer (Eq. (125)).
@@ F.2 Null mean activation function
We now explore the converse scenario, namely an activation function $g\left({\cdot}\right)$ satisfying:
$\left\langle{g^{(1)}_{i}}\right\rangle\equiv\left\langle{g\left({h^{(1)}_{i}}\right)}\right\rangle\stackrel{{\scriptstyle}}{{=}}\left\langle{g^{(1)}}\right\rangle=0.$ (127)
This equality arises because the set $\{h^{(1)}_{i}\}$ is identically distributed. Unlike the non-null mean case ((125)), this scenario leads to:
$p_{h^{(2)}_{i}}^{(\chi)}\left({x}\right)\xrightarrow{N_{1}\rightarrow\infty}\mathcal{N}\left(x;0,\textrm{Var}_{\chi}\left({g^{(1)}}\right)S_{\boldsymbol{w}_{i}^{2}}\right)=\mathcal{N}\left(x;0,\textrm{Var}_{\chi}\left({g^{(1)}}\right)\mkern 1.5mu\overline{\mkern-1.5muS_{\boldsymbol{w}_{i}^{2}}\mkern-1.5mu}\mkern 1.5mu\right).$ (128)
@@ F.2.1 Anti-Symmetric activation functions
Assume $p_{O^{(c)}}^{(\chi)}\left({x}\right)$ asymptotically converges to a Gaussian distribution with non-zero variance (see App. D.3). If $p_{\mu_{c}}\left({x}\right)=\delta\left(x-a\right),\;\forall c\,$, (129) for some $a\in\mathbb{R}$, we have an absence of IGB. Conversely, if condition (129) is not satisfied, IGB emerges, leading to disproportionate values of $\{f_{i}\}$ even in the infinite dataset size limit, thus excluding finite size effects.
@@
$g^{(l)}_{i}=g\left(h^{(l)}_{i}\right)=\tanh\left(h^{(l)}_{i}\right),\;l\in\{0,\dots,L\},\;i\in\{0,\dots,N_{l}\}.$ (130)
Recalling Eq. (62), the asymptotic distribution of $p_{h^{(1)}_{i}}^{(\chi)}\left({x}\right)$ is known and remains independent of the activation function choice. The question arises: ”How is this distribution transformed after passing through the activation function (Eq. (130))?”
@@
$p_{X}\left({x}\right)dx=p_{Y}\left({y(x)}\right)dy,$ (131)
@@
$p_{g^{(1)}_{i}}^{(\chi)}\left({y}\right)=\frac{e^{-\frac{(\operatorname{atanh}(y))^{2}}{2\sigma^{2}_{h^{(1)}}}}}{\sqrt{2\pi\sigma^{2}_{h^{(1)}}}}\frac{1}{(1-y^{2})}.$ (132)
We observe from Eq. (132) that the distribution is symmetric, i.e., $p_{g^{(1)}_{i}}^{(\chi)}\left({y}\right)=p_{g^{(1)}_{i}}^{(\chi)}\left({-y}\right)$. This symmetry arises directly from the anti-symmetric nature of the activation function combined with the symmetry of $p_{h^{(1)}_{i}}^{(\chi)}\left({x}\right)$. This would also hold true for other anti-symmetric activation functions.
@@
$\left\langle{g^{(1)}_{i}}\right\rangle=0\Longrightarrow\left\langle{h^{(2)}_{j}}\right\rangle=0,\;\;\;\forall i,j.$ (133)
This is a stark contrast to the scenarios discussed in the previous section, as it means all nodes $\{h^{(2)}_{j}\}$ are identically distributed. Specifically,
$p_{\left\langle{h^{(2)}_{i}}\right\rangle}\left({x}\right)=\delta\left(x\right)\;\forall i.$ (134)
As we continue to propagate through the network’s remaining hidden layers, this pattern persists. Given the similarity between $p_{h^{(2)}_{i}}^{(\chi)}\left({x}\right)$ and $p_{h^{(1)}_{i}}^{(\chi)}\left({x}\right)$, the same reasoning can be applied layer by layer, leading to
$p_{\left\langle{h^{(3)}_{i}}\right\rangle}\left({x}\right)=\delta\left(x\right)\;\forall i,$ (135)
@@
$p_{\mu_{c}}\left({x}\right)=\delta\left(x\right)\;\forall c.$ (136)
Therefore, by employing an anti-symmetric activation function such as $\tanh$, IGB is not observed, under the assumption that the chosen initialization maintains a non-zero asymptotic variance for $p_{O^{(c)}}^{(\chi)}\left({x}\right)$.
@@ F.3 Eliminate/Trigger IGB with a generic activation function
Figure F.1: Contrast between SReLU and ReLU. Shifting ReLU to meet Eq. (128) conditions eliminates IGB. The figure shows the differing regimes through $p_{f_{0}}\left({x}\right)$. Dataset GB and model SHLP were used for these simulations (see App. I.4 for details).
@@ G.1 Effect of data standardization
$\left\langle{\rho^{(1;1)}_{i}}\right\rangle=\left\langle{h^{(1)}_{i}{}}\right\rangle=\left\langle{\sum_{b}(\xi_{b}^{(a)}+K)w^{(0)}_{ib}}\right\rangle=K\sum_{b}w^{(0)}_{ib}$, (139)
@@
$\left\langle{\rho^{(1;1)}_{i}}\right\rangle=\left\langle{h^{(1)}_{i}{}}\right\rangle\sim\mathcal{N}\left(0,K^{2}\sigma_{w}^{2}\right)\,.$ (140)
@@
$\lim_{N_{1}\rightarrow\infty}p_{\mu_{c}}\left({x}\right)=\mathcal{N}\left(x;0,\sigma_{w}^{4}K^{2}\right)\,\,.$ (142)
@@
$p_{O^{(c)}}^{(\chi)}\left({x}\right)=\mathcal{N}\left(x;\mu_{c},\sum_{j}\left(w^{(1)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({h^{(1)}_{j}}\right)\right)\,$. (143)
@@
$\lim_{d,N_{1}\rightarrow\infty}\textrm{Var}_{\chi}\left({O^{(c)}}\right)=\sigma_{w}^{4}\,.$ (144)
@@
Thus, the ratio $\frac{\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)}{\textrm{Var}_{\chi}\left({O^{(c)}}\right)}$ diverges as $|K|$ increases, indicating a significant amplification of IGB.
@@
Following a similar approach as outlined in App. C.2.1, particularly considering Eq. (LABEL:eq:<h>^l+1) and Eq. (41), in our specific case (i.e., with $h^{(l+1)}_{c}=O^{(c)}$), we obtain:
$\lim_{N_{1}\rightarrow\infty}\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)=\sigma_{w}^{2}\int_{-\infty}^{\infty}\left\langle{g^{(l)}_{j}}\right\rangle^{2}p_{\left\langle{h^{(1)}_{j}{}}\right\rangle}\left({x}\right)dx\,$, (147)
where the integral on the r.h.s. represents the expectation of $\left\langle{g^{(l)}_{j}}\right\rangle^{2}$ over the population $\left\{\left\langle{g^{(1)}_{j}}\right\rangle\right\}$; specifically, as $\left\langle{g^{(l)}_{j}}\right\rangle$ is a function of $\left\langle{h^{(1)}_{i}{}}\right\rangle$, we can express the mean in terms of the measure of the latter. Starting from Eq. (147), we have
$\lim_{N_{1}\rightarrow\infty}\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)\stackrel{{\scriptstyle\boldsymbol{a}}}{{>}}\penalty 10000\ \penalty 10000\$$\sigma_{w}^{2}\int_{0}^{\infty}\left\langle{g^{(l)}_{j}}\right\rangle^{2}p_{\left\langle{h^{(1)}_{j}{}}\right\rangle}\left({x}\right)dx\stackrel{{\scriptstyle\boldsymbol{b}}}{{>}}\sigma_{w}^{2}\int_{0}^{\infty}\left\langle{h^{(1)}_{j}{}}\right\rangle^{2}p_{\left\langle{h^{(1)}_{j}{}}\right\rangle}\left({x}\right)dx\penalty 10000\ \penalty 10000\ \stackrel{{\scriptstyle\boldsymbol{c}}}{{=}}\penalty 10000\ \penalty 10000\$(148)
@@
$\lim_{d,N_{1}\rightarrow\infty}\textrm{Var}_{\chi}\left({O^{(c)}}\right)<\sigma_{w}^{4}\,.$ (150)
@@ G.2 Effect of max pooling
Consider a dataset with Gaussian-distributed, i.i.d. components, where $\xi_{b}^{(a)}\sim\mathcal{N}(0,1)$, and $\xi_{b}^{(a)}$ represents the $b$-th component of the $a$-th input vector. Inputs are processed through an MLP with a single hidden layer ($L=1$), mapping the input to output according to Eq. (21), Eq. (22), and Eq. (23). The weights are initialized following the Kaiming normal scheme, i.e., $w^{(l)}_{ij}\sim\mathcal{N}\left(0,\frac{\sigma_{w}^{2}}{N_{l-1}}\right)$, with zero bias weights. In a setting with max pooling layer and ReLU activation function, an increase in the kernel size, $m$, leads to an asymptotic amplification of the Initial Guessing Bias (IGB). Formally: $\lim_{m\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (152)
@@
Consider the set $\{g^{(1)}_{i}\}_{i=1}^{m}$ such that
$p_{g^{(1)}_{i}}^{(\chi)}\left({x}\right)=\frac{1}{2}\delta\left(x\right)+\Theta\left(x\right)\mathcal{N}\left(x;0,\sigma_{w}^{2}\right),$ (153)
for a fixed set $\mathcal{W}$. Now, define the set of random variables $\{\omega_{i}\}_{i=1}^{m}$, where $n=\log(m)$, as follows:
$\omega_{i}\equiv\frac{g^{(1)}_{i}}{\sqrt{\log(m)}}=\frac{g^{(1)}_{i}}{\sqrt{n}}.$ (154)
@@
$p_{\omega}^{(\chi)}\left({\omega(x)}\right)d\omega=p_{g^{(1)}}^{(\chi)}\left({x}\right)dg^{(1)},$ (155)
@@
$p_{\omega_{i}}^{(\chi)}\left({x}\right)=\frac{1}{2}\delta\left(x\right)+\Theta\left(x\right)\mathcal{N}\left(x;0,\frac{\sigma_{w}^{2}}{n}\right).$ (156)
@@
We have shown that, as $m\to\infty$, $\max\{\omega_{i}\}$ converges to a deterministic value, implying that
$\lim_{m\to\infty}\frac{\sqrt{\textrm{Var}\left({\max\{\omega_{i}\}}\right)}}{\mathbb{E}\left({\max\{\omega\}}\right)}=0.$ (161)
@@
$\lim_{m\to\infty}\frac{\sqrt{\textrm{Var}_{\chi}\left({\rho^{(1;m)}}\right)}}{\left\langle{\rho^{(1;m)}}\right\rangle}=\frac{\cancel{\log(m)}\sqrt{\textrm{Var}\left({\max\{\omega\}}\right)}}{\cancel{\log(m)}\mathbb{E}\left({\max\{\omega\}}\right)}=0.$ (162)
@@
$\lim_{m\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\frac{\left\langle{\rho^{(1;m)}}\right\rangle^{2}}{\textrm{Var}_{\chi}\left({\rho^{(1;m)}}\right)}=\infty.$ (163)
In the proof of Thm. G.2, we focused on the statistical behavior of the maximum value in a growing set of random variables (r.v.s), particularly emphasizing positive event occurrences. This analysis remains applicable if we substitute the ReLU activation function with a linear one, as ReLU does not change the distribution within the positive domain of its support. This implies that the amplification of Initial Guessing Bias (IGB) in models with max pooling is independent of the choice of activation function, manifesting even without any activation function. This finding is not surprising, given that App. D.1 demonstrates that the key factor in the emergence of IGB is the statistics of nodes post-pooling layer, in particular $\left\langle{\rho^{(1;m)}}\right\rangle$. Hence, when considering models with pooling layers, it’s crucial to evaluate the joint impact of the activation function and the pooling layer to accurately assess the presence and intensity of IGB.

### v5

@@ F.1 Non-null mean activation function
Let us consider a generic activation function $H\left({\cdot}\right)$, such that
$\left\langle{H^{(1)}_{i}}\right\rangle\equiv\left\langle{H\left({Z^{(1)}_{i}}\right)}\right\rangle\stackrel{{\scriptstyle}}{{=}}\left\langle{H^{(1)}}\right\rangle\neq 0.$ (124)
where the second step follow from the fact that $\{Z^{(1)}_{i}\}$ are identically distributed. Starting from this hypothesis on the activation function (Eq. (124)) we can follow exactly the same analysis presented in App. G.3.1. In particular:
@@
$f_{O_{c}}^{(\chi)}\left({o}\right)=f_{Z^{(2)}_{c}}^{(\chi)}\left({o}\right)\,.$ (126)
For a generic deep architecture we can keep following the analysis of App. G.3.1, substituiting the right expression to $f_{H^{(2)}_{i}}^{(\chi)}\left({h{}{}}\right)$ dependent to the specific used activation function. Proceeding in this way we will find, for each layer, a symmetry breaking in the nodes distribution, similarly to what we observed in the second layer (Eq. (125)).
@@ F.2 Null mean activation function
We now explore the converse scenario, namely an activation function $H\left({\cdot}\right)$ satisfying:
$\left\langle{H^{(1)}_{i}}\right\rangle\equiv\left\langle{H\left({Z^{(1)}_{i}}\right)}\right\rangle\stackrel{{\scriptstyle}}{{=}}\left\langle{H^{(1)}}\right\rangle=0.$ (127)
This equality arises because the set $\{Z^{(1)}_{i}\}$ is identically distributed. Unlike the non-null mean case ((125)), this scenario leads to:
$f_{Z^{(2)}_{i}}^{(\chi)}\left({z{}{}}\right)\xrightarrow{N_{1}\rightarrow\infty}\mathcal{N}\left(z{}{};0,\textrm{Var}_{\chi}\left({H^{(1)}}\right)S_{\boldsymbol{w}_{i}^{2}}\right)=\mathcal{N}\left(z{}{};0,\textrm{Var}_{\chi}\left({H^{(1)}}\right)\mkern 1.5mu\overline{\mkern-1.5muS_{\boldsymbol{w}_{i}^{2}}\mkern-1.5mu}\mkern 1.5mu\right).$ (128)
@@ F.2.1 Anti-Symmetric activation functions
Assume $f_{O_{c}}^{(\chi)}\left({o}\right)$ asymptotically converges to a Gaussian distribution with non-zero variance (see App. D.3). If $f_{\mu_{c}}\left({m}\right)=\delta\left(m-a\right),\;\forall c\,$, (129) for some $a\in\mathbb{R}$, we have an absence of IGB. Conversely, if condition (129) is not satisfied, IGB emerges, leading to disproportionate values of $\{g_{i}\}$ even in the infinite dataset size limit, thus excluding finite size effects.
@@
$H^{(l)}_{i}=g\left(Z^{(l)}_{i}\right)=\tanh\left(Z^{(l)}_{i}\right),\;l\in\{0,\dots,L\},\;i\in\{0,\dots,N_{l}\}.$ (130)
Recalling Eq. (62), the asymptotic distribution of $f_{Z^{(1)}_{i}}^{(\chi)}\left({z{}{}}\right)$ is known and remains independent of the activation function choice. The question arises: ”How is this distribution transformed after passing through the activation function (Eq. (130))?”
@@
$f_{X}\left({x}\right)dx=f_{Y}\left({y(x)}\right)dy,$ (131)
@@
$f_{H^{(1)}_{i}}^{(\chi)}\left({h{}{}}\right)=\frac{e^{-\frac{(\operatorname{atanh}(h{}{}))^{2}}{2\sigma^{2}_{Z^{(1)}}}}}{\sqrt{2\pi\sigma^{2}_{Z^{(1)}}}}\frac{1}{(1-h{}{}^{2})}.$ (132)
We observe from Eq. (132) that the distribution is symmetric, i.e., $f_{H^{(1)}_{i}}^{(\chi)}\left({h{}{}}\right)=f_{H^{(1)}_{i}}^{(\chi)}\left({-h{}{}}\right)$. This symmetry arises directly from the anti-symmetric nature of the activation function combined with the symmetry of $f_{Z^{(1)}_{i}}^{(\chi)}\left({z{}{}}\right)$. This would also hold true for other anti-symmetric activation functions.
@@
$\left\langle{H^{(1)}_{i}}\right\rangle=0\Longrightarrow\left\langle{Z^{(2)}_{j}}\right\rangle=0,\;\;\;\forall i,j.$ (133)
This is a stark contrast to the scenarios discussed in the previous section, as it means all nodes $\{Z^{(2)}_{j}\}$ are identically distributed. Specifically,
$f_{\left\langle{Z^{(2)}_{i}}\right\rangle}\left({m_{z{}{}}}\right)=\delta\left(m_{z{}{}}\right)\;\forall i.$ (134)
As we continue to propagate through the network’s remaining hidden layers, this pattern persists. Given the similarity between $f_{Z^{(2)}_{i}}^{(\chi)}\left({z{}{}}\right)$ and $f_{Z^{(1)}_{i}}^{(\chi)}\left({z{}{}}\right)$, the same reasoning can be applied layer by layer, leading to
$f_{\left\langle{Z^{(3)}_{i}}\right\rangle}\left({m_{z{}{}}}\right)=\delta\left(m_{z{}{}}\right)\;\forall i,$ (135)
@@
$f_{\mu_{c}}\left({m}\right)=\delta\left(m\right)\;\forall c.$ (136)
Therefore, by employing an anti-symmetric activation function such as $\tanh$, IGB is not observed, under the assumption that the chosen initialization maintains a non-zero asymptotic variance for $f_{O_{c}}^{(\chi)}\left({o}\right)$.
@@ F.3 Eliminate/Trigger IGB with a generic activation function
Figure F.1: Contrast between SReLU and ReLU. Shifting ReLU to meet Eq. (128) conditions eliminates IGB. The figure shows the differing regimes through $f_{G_{0}}\left({g_{0}}\right)$. Dataset GB and model SHLP were used for these simulations (see App. I.4 for details).
@@ G.1 Effect of data standardization
$\left\langle{U^{(1;1)}_{i}}\right\rangle=\left\langle{Z^{(1)}_{i}{}}\right\rangle=\left\langle{\sum_{b}(\xi_{b}^{(a)}+K)w^{(0)}_{ib}}\right\rangle=K\sum_{b}w^{(0)}_{ib}$, (139)
@@
$\left\langle{U^{(1;1)}_{i}}\right\rangle=\left\langle{Z^{(1)}_{i}{}}\right\rangle\sim\mathcal{N}\left(0,K^{2}\sigma_{w}^{2}\right)\,.$ (140)
@@
$\lim_{N_{1}\rightarrow\infty}f_{\mu_{c}}\left({m}\right)=\mathcal{N}\left(m;0,\sigma_{w}^{4}K^{2}\right)\,\,.$ (142)
@@
$f_{O_{c}}^{(\chi)}\left({o}\right)=\mathcal{N}\left(o;\mu_{c},\sum_{j}\left(w^{(1)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({Z^{(1)}_{j}}\right)\right)\,$. (143)
@@
$\lim_{d,N_{1}\rightarrow\infty}\textrm{Var}_{\chi}\left({O_{c}}\right)=\sigma_{w}^{4}\,.$ (144)
@@
Thus, the ratio $\frac{\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)}{\textrm{Var}_{\chi}\left({O_{c}}\right)}$ diverges as $|K|$ increases, indicating a significant amplification of IGB.
@@
Following a similar approach as outlined in App. C.2.1, particularly considering Eq. (LABEL:eq:<h>^l+1) and Eq. (41), in our specific case (i.e., with $Z^{(l+1)}_{c}=O_{c}$), we obtain:
$\lim_{N_{1}\rightarrow\infty}\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)=\sigma_{w}^{2}\int_{-\infty}^{\infty}\left\langle{H^{(l)}_{j}}\right\rangle^{2}f_{\left\langle{Z^{(1)}_{j}{}}\right\rangle}\left({m_{z{}{}}}\right)dm_{z{}{}}\,$, (147)
where the integral on the r.h.s. represents the expectation of $\left\langle{H^{(l)}_{j}}\right\rangle^{2}$ over the population $\left\{\left\langle{H^{(1)}_{j}}\right\rangle\right\}$; specifically, as $\left\langle{H^{(l)}_{j}}\right\rangle$ is a function of $\left\langle{Z^{(1)}_{i}{}}\right\rangle$, we can express the mean in terms of the measure of the latter. Starting from Eq. (147), we have
$\lim_{N_{1}\rightarrow\infty}\textrm{Var}_{\mathcal{W}}\left({\mu_{c}}\right)\stackrel{{\scriptstyle\boldsymbol{a}}}{{>}}\penalty 10000\ \penalty 10000\$$\sigma_{w}^{2}\int_{0}^{\infty}\left\langle{H^{(l)}_{j}}\right\rangle^{2}f_{\left\langle{Z^{(1)}_{j}{}}\right\rangle}\left({m_{z{}{}}}\right)dm_{z{}{}}\stackrel{{\scriptstyle\boldsymbol{b}}}{{>}}\sigma_{w}^{2}\int_{0}^{\infty}\left\langle{Z^{(1)}_{j}{}}\right\rangle^{2}f_{\left\langle{Z^{(1)}_{j}{}}\right\rangle}\left({m_{z{}{}}}\right)dm_{z{}{}}\penalty 10000\ \penalty 10000\ \stackrel{{\scriptstyle\boldsymbol{c}}}{{=}}\penalty 10000\ \penalty 10000\$(148)
@@
$\lim_{d,N_{1}\rightarrow\infty}\textrm{Var}_{\chi}\left({O_{c}}\right)<\sigma_{w}^{4}\,.$ (150)
@@ G.2 Effect of max pooling
Consider a dataset with Gaussian-distributed, i.i.d. components, where $\xi_{b}^{(a)}\sim\mathcal{N}(0,1)$, and $\xi_{b}^{(a)}$ represents the $b$-th component of the $a$-th input vector. Inputs are processed through an MLP with a single hidden layer ($L=1$), mapping the input to output according to Eq. (21), Eq. (22), and Eq. (23). The weights are initialized following the Kaiming normal scheme, i.e., $w^{(l)}_{ij}\sim\mathcal{N}\left(0,\frac{\sigma_{w}^{2}}{N_{l-1}}\right)$, with zero bias weights. In a setting with max pooling layer and ReLU activation function, an increase in the kernel size, $k$, leads to an asymptotic amplification of the Initial Guessing Bias (IGB). Formally: $\lim_{m\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\infty.$ (152)
@@
Consider the set $\{H^{(1)}_{i}\}_{i=1}^{k}$ such that
$f_{H^{(1)}_{i}}^{(\chi)}\left({h{}{}}\right)=\frac{1}{2}\delta\left(h{}{}\right)+\Theta\left(h{}{}\right)\mathcal{N}\left(h{}{};0,\sigma_{w}^{2}\right),$ (153)
for a fixed set $\mathcal{W}$. Now, define the set of random variables $\{\omega_{i}\}_{i=1}^{k}$, where $n=\log(k)$, as follows:
$\omega_{i}\equiv\frac{H^{(1)}_{i}}{\sqrt{\log(k)}}=\frac{H^{(1)}_{i}}{\sqrt{n}}.$ (154)
@@
$f_{\omega_{i}}^{(\chi)}\left({\omega}\right)d\omega=f_{H^{(1)}}^{(\chi)}\left({h{}{}(\omega)}\right)dh{}{},$ (155)
@@
$f_{\omega_{i}}^{(\chi)}\left({\omega}\right)=\frac{1}{2}\delta\left(\omega\right)+\Theta\left(\omega\right)\mathcal{N}\left(\omega;0,\frac{\sigma_{w}^{2}}{n}\right).$ (156)
@@
We have shown that, as $k\to\infty$, $\max\{\omega_{i}\}$ converges to a deterministic value, implying that
$\lim_{k\to\infty}\frac{\sqrt{\textrm{Var}\left({\max\{\omega_{i}\}}\right)}}{\mathbb{E}\left({\max\{\omega\}}\right)}=0.$ (161)
@@
$\lim_{k\to\infty}\frac{\sqrt{\textrm{Var}_{\chi}\left({U^{(1;k)}}\right)}}{\left\langle{U^{(1;k)}}\right\rangle}=\frac{\cancel{\log(k)}\sqrt{\textrm{Var}\left({\max\{\omega\}}\right)}}{\cancel{\log(k)}\mathbb{E}\left({\max\{\omega\}}\right)}=0.$ (162)
@@
$\lim_{k\rightarrow\infty}\gamma(\mathcal{A},\psi(\chi))=\frac{\left\langle{U^{(1;k)}}\right\rangle^{2}}{\textrm{Var}_{\chi}\left({U^{(1;k)}}\right)}=\infty.$ (163)
In the proof of Thm. G.2, we focused on the statistical behavior of the maximum value in a growing set of random variables (r.v.s), particularly emphasizing positive event occurrences. This analysis remains applicable if we substitute the ReLU activation function with a linear one, as ReLU does not change the distribution within the positive domain of its support. This implies that the amplification of Initial Guessing Bias (IGB) in models with max pooling is independent of the choice of activation function, manifesting even without any activation function. This finding is not surprising, given that App. D.1 demonstrates that the key factor in the emergence of IGB is the statistics of nodes post-pooling layer, in particular $\left\langle{U^{(1;k)}}\right\rangle$. Hence, when considering models with pooling layers, it’s crucial to evaluate the joint impact of the activation function and the pooling layer to accurately assess the presence and intensity of IGB.